if a and d are two complex numbers ,then the sum to (n+1) terms of the following series
aC0-(a+d)C1+(a+2d)C2-......... is

Grade:11

2 Answers

Chetan Mandayam Nayakar
12 years ago

splitting up the series to two series, we get

a(C0-C1+C2...(-1)nCn)+d(-C1+2C2...(-1)nnCn)=a(1-x)n+d(d((1-x)n)/dx, where x is equal to unity,substituting, we get

 sum of each of the two series is zero. Therefore the answer is zero.

Dear moderator: If more details are required, please contact me at mn_chetan@yahoo.com

Ashwin Muralidharan IIT Madras
12 years ago

Hi Himanshu,

 

Take a common and you have a{C0 - C1 + C2 -.......}     -------------- (1)

Take -b common and you have -b{C1 - 2C2 + 3C3 -........}  ------------- (2)

 

In the expansion of (1-x)^n, put x=1, you will get (1) = 0

And for the terms in bracket of (2), differentiate (1-x)^n, and put x=1

ie put x=1 in -n*(1-x)^n-1, which is again 0.

 

And hence the answer is 0.

 

Best Regards,

Ashwin (IIT Madras).